CBT has reported EBIT of $100mn this year. Its total capital expenditure is $20mn and depreciation is $10mn this year. This year’s account receivables and account payables are $5mn and $2mn, respectively. The net working capital was $10mn last year and $20mn this year, respectively. Both the company’s EBIT and investment (including tangible and working capital) needs are expected to grow at a constant rate of 1% per year. It is expected that CBT maintains the current debt-to-equity ratio of 1. The corporate tax rate is 40%. The required return on equity is 17%. The cost of the debt capital is 5%. The number of outstanding shares is 100mn. Its net debt is $250mn. Obtain the value of a share based on the discounted free cash flow model. Explain and justify your procedure in detail.
Answer #1

Current Free cash flows to the firm = C0 = EBIT x (1 - T) + Depreciation - Increase in working capital - Capital expenditure = 100 x (1 - 40%) + 10 - (20 - 10) - 20 = $ 40 million

Expected free cash flow = C1 = C0 x (1 + g) = 40 x (1 + 1%) = $ 40.40 mn

D/E = 1; hence Wd = D / (D + E) = 1 / (1 + 1) = 0.5 and We = 1 - Wd = 1 - 0.5 = 0.5

WACC = r = Wd x Kd x (1 - T) + We x Ke = 0.5 x 5% x (1 - 40%) + 0.5 x 17% = 10%

Hence, value of the operations = C1 / (r - g) = 40.40 / (10% - 1%) = $ 448.89 million

Equity value = Value of the operations - Net debt = 448.89 - 250 = $ 198.89 million

The fair value of the share = Equity value / N = 198.89 / 100 = $ 1.9889 = $ 1.99 / share